PLEASE HELP ME SOLVE THIS QUESTION...
1/9 = 27^4X-3

Answer by GReddy(2)

Solutions: X=5.85410442760553e-06.
Graphical form: simplifies to
Text form: 1/9=27^4X-3 simplifies to 0=0
Cartoon (animation) form:
simplify_cartoon( 1/9=27^4X-3 )
--------------------------------------------------------------------------------
Detailed explanation:
--------------------------------------------------------------------------------
Look at .
Replace constants in factor: , with
It becomes .
--------------------------------------------------------------------------------
Look at .
Moved these terms to the left ,
It becomes .
--------------------------------------------------------------------------------
Look at .
Added fractions or integers together
It becomes .
--------------------------------------------------------------------------------
Look at .
Moved to the right of expression
It becomes .
--------------------------------------------------------------------------------
Look at .
Solved linear equation equivalent to -531441*X+3.11111111111111 =0
It becomes .
Result:
Solutions: X=5.85410442760553e-06.
Done!